# A Prospect-Theoretic Policy Gradient Framework for Behaviorally Nuanced Reinforcement Learning

### v3

@@ G.3 CPT value examples
[…]
Table 1: CPT value examples. *: Note that $w^{+}$ and $w^{-}$ are discontinuous for VaR and CVaR.
G.4 Proof: CVaR, Var and distortion risk measures are CPT values
Proposition 9.
Any distortion risk measure of a given random variable $X$ can be written as a CPT value with $u^{+}=\text{id}^{+}$, $u^{-}=-\text{id}^{-}$, $w^{+}=\tilde{g}$ and $w^{-}=g\,.$
Proof.
We recognize the CPT-value of $X$ with $u^{+}=\text{id}^{+}$, $u^{-}=-\text{id}^{-}$, $w^{+}=\tilde{g}$ and $w^{-}=g\,.$
Remark 10.
G.5 Simple examples and further insights about CPT
Therefore, due to its failure to capture such settings as a descriptive model, the standard expected utility theory has been called into question by the pioneering behavioral psychologist Daniel Kahneman together with his colleague Amos Tversky [Kahneman and Tversky, 1979]. In particular, Daniel Kahneman has been awarded the Nobel Prize in Economic Sciences in 2002 ”for having integrated insights from psychological research into economic science, especially concerning human judgment and decision-making under uncertainty”. In their seminal works combining cognitive psychology and economics, they laid the foundations of the so-called prospect theory and its cumulative version later on [Tversky and Kahneman, 1992] to explain several empirical observations invalidating the standard expected utility theory.
G.6 Connection to General Utility RL and Convex RL in finite trials
$\max_{\pi}F(d_{\rho}^{\pi})\,$, (2)
where $F$ is the real valued utility function defined on the set of probability measures over the state or state-action space, $\rho$ is the initial state distribution and $d_{\rho}^{\pi}$ is the state (or sometimes state-action) occupancy measure induced by the policy $\pi\,.$ This problem captures the standard RL problem as a particular case by considering a linear functional $F$ defined using a fixed given reward function. Recently, motivated by practical concerns, Mutti et al. [2023b] argued for the relevance of a variation of the problem under the qualification of convex RL in finite trials. They introduce for this the empirical state distributions $d_{n}\in\Delta(\mathcal{S})$ defined for every state $s\in\mathcal{S}$ by:
$\max_{\pi}\xi_{n}(\pi):=\mathbb{E}[F(d_{n}^{\pi})]\,.$ (4)
Note that $d_{n}^{\pi}$ is a random variable as it is an empirical state distribution. Observe also that $\lim_{n\to\infty}\xi_{n}(\pi)=F(d_{\rho}^{\pi})$ under mild technical conditions (e.g. continuity and boundedness of $F$). This shows the connection between the above final trial convex RL objective and the general utility RL problem (2). The interesting differences between both problem formulations arise for small values of $n$. Of particular interest, both in this paper and in Mutti et al. [2023b], is the single trial RL setting where $n=1$.
which is of the form $\xi_{1}(\pi)$, the single-trial RL objective as defined in Mutti et al. [2023b]. Indeed, it suffices to write the following to observe it:
G.7 Choice of the utility function
Appendix H Proofs for Section 2.3
H.1 The need for stochastic policies
To prove the result (i.e. the need for stochastic policies), we consider a simple MDP with only two states (an initial state and a terminal one) and two actions (A and B). See Fig. 14(a) below. We choose the identity as utility. Action A yields reward 1 with probability 1 and action B yields either 0 or $\frac{3}{2}$ with probability $\frac{1}{2}$ each. We further consider the following probability distortion function $w^{+}:[0,1]\to[0,1]$ defined for every $x\in[0,1]$ as follows:
[…]
and we set $w^{-}=0$. All the policies can be described with a single scalar $p\in[0,1]$, the probability of choosing B instead of A.
@@ H.1 The need for stochastic policies
$\mathbb{C}(X)=w^{+}\left(1-\frac{p}{2}\right)+\frac{1}{2}w^{+}\left(\frac{p}{2}\right)\,.$ (8)
@@
  • For the policy corresponding to $p=1$, $\mathbb{C}(X)=\frac{3}{2}w^{+}(\frac{1}{2})=\frac{13}{12}\approx 1.08\,.$
@@
$\mathbb{C}(X)=w^{+}(0.9)+\frac{1}{2}w^{+}(0.1)=\frac{17}{18}+\frac{1}{4}=\frac{43}{36}\approx 1.19$
@@ H.2 Proof of Proposition 2
$w^{+}(x)=\begin{cases}5x&\text{if }x\leq 0.1,\\ \frac{1}{2}+\frac{5}{9}(x-0.1)&\text{otherwise}.\end{cases}$
We also set $w^{-}=0$. However, we consider another MDP. Our MDP has three states: an initial state $s_{0}$, an intermediate state $s_{1}$, and a terminal state $s_{2}$. There are two actions: A and B. All trajectories start in $s_{0}$. Any action from $s_{0}$ leads to $s_{1}$ with probability $1$ and yields reward +1+1 with probability $\frac{1}{2}$ and 0 otherwise. The action taken when in $s_{0}$ is completely irrelevant. Any action taken in $s_{1}$ leads to $s_{2}$ with probability $1$ and the episode stops as soon as $s_{2}$ is reached. When taking action A in $s_{1}$, the reward is either 0 or +2+2, with probability $\frac{1}{2}$ each. When taking action B in $s_{1}$, the reward is +1+1 with probability $1$. All policies in $\Pi_{NM}$ can be described by $(p_{\text{start}},p_{0},p_{1})$, where $p_{\text{start}}$ is the probability of choosing action A when in $s_{0}$, $p_{0}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward 0 and $p_{1}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward $1$. $p_{\text{start}}$ is irrelevant to the performance of the policy so we can ignore it. The set of Markovian policies here is the set of policies such as $p_{0}=p_{1}$. $\mathbb{C}(\pi)$ is a piecewise affine function of $p_{0}$ and $p_{1}$ and it can therefore be directly maximized. We omit the calculations here: one can check that the best achievable CPT value for Markovian policies is $\approx 0.616$ for $p_{0}=p_{1}=0.4$ but that a CPT value of $\approx 0.625$ is achievable for $p_{0}=0$ and $p_{1}=0.4$, proving the lemma.
Appendix I Proofs and Additional Details for Section 3
I.1 Proof of Theorem 3
@@
I.2 Alternative Practical Procedure for Computing Policy Gradients
@@
Appendix J Proofs for Section 4
@@
J.1 Proof of Proposition 5: Consistency of the PG estimator
@@
J.2 Proof of Proposition 6: Sample complexity of the PG estimator
@@
We start with the random variable $Y_{n}$ and show how to control $\hat{\phi}_{n}^{+}$, $\hat{\phi}_{n}^{-}$ can be bounded in the same way. Recalling the rewritten form of $\hat{\phi}_{n}^{+}$ in (J.1), it appears that $\hat{\phi}_{n}^{+}$ is clearly comparable to $\varphi^{+}(v):=\int_{z=0}^{\max(v,0)}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z))dz$ as defined in Theorem 3.
The rest of the proof for controlling this quantity follows similar lines to the proof of Proposition 3 in L.A. et al. [2016]. Since $u^{+}(\tau)$ is supposed to be bounded by $M_{u}$ and $w_{+}^{\prime}$ (note here this is the derivative of $w_{+}$ unlike in [L.A. et al., 2016, Proposition 3]) is $L_{w}$-Lipschitz, we have
@@
J.3 Proof of Proposition 7: Asymptotic convergence
This result follows from a standard application of the stochastic approximation framework (see e.g. Borkar [2008], Benaïm [2006]) similarly to L.A. et al. [2016]. Under our regularity assumptions on the utility and weight functions, we can show asymptotic convergence of the iterates to the set of stationary points of our CPT objective. The idea is to show that the iterates track a gradient flow defined by the gradient of the CPT-PO objective with vanishing step sizes. The proof follows the same steps as the proof of Theorem 1 in L.A. et al. [2016] (see appendix D, p. 25), so we do not reproduce it here, up to the difference that our policy gradients are not estimated using zeroth-order information but rather using our PG Thm. 3 and Prop. 4.
@@
Appendix K More Details about Section 5 and Additional Experiments
@@
Hardware and execution time. We have run part of the experiments on a laptop with a 13th Gen Intel Core i7-1360P2.20 GHz CPU and 32 GB of RAM, and the remaining simulations (section 5, App. K.8, K.9) on a MacBook Pro M4 with 48 GB of RAM. Experiments can take a few seconds for the simplest ones, to a few minutes for the ones in the main part. More complex ones take a couple of hours, e.g. for electricity management (App. K.7), Finance (App. K.8) and MuJoCo (App. K.9).
@@
Weight functions. The risk-neutral $w^{+}$ function is simply the identity function. As for the definition of other probability distortion functions $w^{+}$ we use for experiments, we define:
@@
K.1 More details about bandit simulations in Section 5
@@
K.2 CPT compared to Distortion Risk Measures
Figure 13: Illustration of the flexibility of CPT compared to the Distortion Risk Measure. Notice how $w^{-}$ is distinct from both $w^{+}$ and $p\mapsto 1-w^{+}(1-p)$.
K.3 Illustration of the need for stochastic policies in CPT-RL
@@
Figure 15: Results on non-deterministic policies and batch size influence, over $100$ runs. The black dots are the medians and the shaded area represents the interquantile range.
@@
K.4 Markovian vs Non-Markovian Policies for CPT-RL
@@
K.6 Traffic Control over a Grid
We simulate a car agent navigating a city grid, where central roads are faster but risk higher delays (see fig. 20). The agent must balance speed against risk by avoiding the city center. In risk-neutral settings, the agent favors faster routes, while risk-averse policies avoid the risky central roads. Fig. 20 (center) demonstrates that our algorithm successfully adapts to different risk-weighted objectives. We also consider solving CPT-PO with different utility functions: risk-neutral identity utility, risk-averse KT utility, as well as exponential utility function. Fig. 23 show the corresponding different CPT returns observed. The obtained policies differ depending on the utility function. For examples of risk-neutral/averse policies obtained, see Fig. 19(b) in appendix K.5.
@@
Figure 21: The probability distortion function $w^{+}$ used for the traffic control experiment.
@@ 2 $\uparrow$ ? $\uparrow$
Figure 23: Returns along the iterations of our PG algorithm for CPT-PO for traffic control. Shaded areas indicate a range of $\pm$ one standard deviation over 20 runs. See Appendix K for details.
Implementation details. In both cases, the risk-neutral optimal solution (going around the city center) is also a local optimum for the risk-averse objective, and, because it is a shorter path, is easier to stumble upon by chance when exploring the MDP. This means we have to implement special measures to force exploration. The algorithm used as is is prone to get stuck from time to time in local minima on this example. It would seem that our $w$ function, which is aggressively risk-averse, hinders exploration. To mitigate this, we introduce an entropy regularization term that we add to the score function with a decaying regularization weight in the policy gradient found in Theorem 3, see App. K for further details. We incorporate entropy regularization in the policy gradient as follows:
@@
K.8 Trading in Financial Markets
@@
  • State space: We consider a seven dimensional continuous state space. Features are constructed from the raw stock market data as previously explained. State transitions are described using the provided time series. See the publicly available code of the environment for more details.
@@
K.9 Control on MuJoCo Environments

### v4

@@ C.1 The need for stochastic policies
and we set $w_{-}=0$. All the policies can be described with a single scalar $p\in[0,1]$, the probability of choosing B instead of A.
@@
$\mathbb{C}(X)=w_{+}\left(1-\frac{p}{2}\right)+\frac{1}{2}w_{+}\left(\frac{p}{2}\right)\,.$ (4)
@@
  • For the policy corresponding to $p=1$, $\mathbb{C}(X)=\frac{3}{2}w_{+}(\frac{1}{2})=\frac{13}{12}\approx 1.08\,.$
@@
$\mathbb{C}(X)=w_{+}(0.9)+\frac{1}{2}w_{+}(0.1)=\frac{17}{18}+\frac{1}{4}=\frac{43}{36}\approx 1.19$
@@ C.2 Proof of Proposition 2
$w_{+}(x)=\begin{cases}5x&\text{if }x\leq 0.1,\\ \frac{1}{2}+\frac{5}{9}(x-0.1)&\text{otherwise}.\end{cases}$
We also set $w_{-}=0$. However, we consider another MDP. Our MDP has three states: an initial state $s_{0}$, an intermediate state $s_{1}$, and a terminal state $s_{2}$. There are two actions: A and B. All trajectories start in $s_{0}$. Any action from $s_{0}$ leads to $s_{1}$ with probability $1$ and yields reward +1+1 with probability $\frac{1}{2}$ and 0 otherwise. The action taken when in $s_{0}$ is completely irrelevant. Any action taken in $s_{1}$ leads to $s_{2}$ with probability $1$ and the episode stops as soon as $s_{2}$ is reached. When taking action A in $s_{1}$, the reward is either 0 or +2+2, with probability $\frac{1}{2}$ each. When taking action B in $s_{1}$, the reward is +1+1 with probability $1$. All policies in $\Pi_{NM}$ can be described by $(p_{\text{start}},p_{0},p_{1})$, where $p_{\text{start}}$ is the probability of choosing action A when in $s_{0}$, $p_{0}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward 0 and $p_{1}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward $1$. $p_{\text{start}}$ is irrelevant to the performance of the policy so we can ignore it. The set of Markovian policies here is the set of policies such as $p_{0}=p_{1}$. $\mathbb{C}(\pi)$ is a piecewise affine function of $p_{0}$ and $p_{1}$ and it can therefore be directly maximized. We omit the calculations here: one can check that the best achievable CPT value for Markovian policies is $\approx 0.616$ for $p_{0}=p_{1}=0.4$ but that a CPT value of $\approx 0.625$ is achievable for $p_{0}=0$ and $p_{1}=0.4$, proving the lemma.
Appendix D Proofs and Additional Details for Section 3
D.1 Proof of Theorem 3
@@
D.2 Alternative Practical Procedure for Computing Stochastic Policy Gradients
@@
Appendix E Proofs for Section 4
@@
E.1 Proof of Proposition 5: Consistency of the PG estimator
@@
E.2 Proof of Proposition 6: Sample complexity of the PG estimator
@@
We start with the random variable $Y_{n}$ and show how to control $\hat{\phi}_{n}^{+}$, $\hat{\phi}_{n}^{-}$ can be bounded in the same way. Recalling the rewritten form of $\hat{\phi}_{n}^{+}$ in (E.1), it appears that $\hat{\phi}_{n}^{+}$ is clearly comparable to $\varphi^{+}(v):=\int_{z=0}^{\max(v,0)}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z))dz$ as defined in Theorem 3.
The rest of the proof for controlling this quantity follows similar lines to the proof of Proposition 3 in L.A. et al. [2016]. Since $u^{+}(\tau)$ is supposed to be bounded by $M_{u}$ and $w_{+}^{\prime}$ (note here this is the derivative of $w_{+}$ unlike in Proposition 3 in L.A. et al. [2016]) is $L_{w}$-Lipschitz, we have
@@
E.3 Proof of Proposition 7: Asymptotic convergence
This result follows from a standard application of the stochastic approximation framework (see e.g. Borkar [2008], Benaïm [2006]) similarly to L.A. et al. [2016]. Under our regularity assumptions on the utility and weight functions, we can show asymptotic convergence of the iterates to the set of stationary points of our CPT objective. The idea is to show that the iterates track a gradient flow defined by the gradient of the CPT-PO objective with vanishing step sizes. The proof follows the same steps as the proof of Theorem 1 in L.A. et al. [2016] (see App. D, p. 25 therein), so we do not reproduce it here, up to the difference that our policy gradients are not estimated using zeroth-order information but rather using our PG Thm. 3 and Prop. 4.
@@ E.3 Proof of Proposition 7: Asymptotic convergence
Appendix F CPT-RL vs Risk-sensitive RL: An Illustration
  • Actions: The agent can move up, down, left, or right.
Results. See figures 8 and 9 and their captions in the next page.
Appendix G Applications of CPT: From Prior Work in Stateless Settings to CPT-RL
G.1 Energy: Renovation and Home Energy Management
G.2 Security: Building Evacuation
G.3 Urban Planning and Mobility
G.4 Finance
G.5 Example: Personalized Treatment for Pain Management
G.6 Further Applications of CPT-RL
  • Electricity management. Please see simulations in the main part (section 5 and App. J.7) in a simple example setting to illustrate our methodology.
  • Finance: portfolio optimization, risk management, behavioral asset pricing (e.g. influence of investor sentiment on price dynamics via e.g. over-weighting of low-probability events, including their preferences). For recent applications of CPT to finance, we refer the reader to a recent paper [Luxenberg et al., 2024] using CPT for portfolio optimization (in a stateless static setting). We also applied our methodology to financial trading (see App. J.8).
Appendix H CPT-RL and Trajectory-Based Reward RL as Preference Learning Paradigms
  1. (a)
  2. (b)
  3. (c)
  4. (d)
Appendix I More About CPT Values and CPT Policy Optimization
I.1 More about Optimal Policies in CPT-RL
The need for stochastic policies. We start our discussion by pointing out a stark difference between optimal policies in standard MDPs and CPT-PO. While there exists an optimal deterministic stationary policy for MDPs (see e.g. Thm. 6.2.10 in Puterman [2014]), this is not the case in general for CPT-PO. Indeed, there does not always exist an optimal policy for CPT-PO in $\Pi_{NM}^{D}\,.$ In general, stochasticity of the policy is essential in solving our CPT-RL problem. To see this, consider an example built around a function $w_{+}$ that puts special emphasis on the $10\%$ of the best outcomes. In this case, the optimal policy needs to be randomized to take advantage of this and obtain the highest returns with some probability without suffering from bad outcomes by deterministically committing to this riskier strategy (see App. C.1 for a proof).
I.2 Positioning CPT-RL in the literature
Remark 19.
I.3 CPT value examples
Setting Utility function $_{+} 𝒘w_{-}$
CPT Any Any Any
EUT Any Identity function Identity function
Distortion risk measure Identity function Any $1-w_{+}(1-t)$
[…]
Table 1: CPT value examples. *: Note that $w_{+}$ and $w_{-}$ are discontinuous for VaR and CVaR.
I.4 Proof: CVaR, Var and distortion risk measures are CPT values
Proposition 20.
Any distortion risk measure of a given random variable $X$ can be written as a CPT value with $u^{+}=\text{id}^{+}$, $u^{-}=-\text{id}^{-}$, $w_{+}=\tilde{g}$ and $w_{-}=g\,.$
Proof.
We recognize the CPT-value of $X$ with $u^{+}=\text{id}^{+}$, $u^{-}=-\text{id}^{-}$, $w_{+}=\tilde{g}$ and $w_{-}=g\,.$
Remark 21.
I.5 Simple examples and further insights about CPT
Therefore, due to its failure to capture such settings as a descriptive model, the standard expected utility theory has been called into question by the pioneering behavioral psychologist Daniel Kahneman together with his colleague Amos Tversky [Kahneman and Tversky, 1979]. In particular, Daniel Kahneman has been awarded the Nobel Prize in Economic Sciences in 2002 ”for having integrated insights from psychological research into economic science, especially concerning human judgment and decision-making under uncertainty”. In their seminal works combining cognitive psychology and economics, they laid the foundations of the so-called prospect theory and its cumulative version later on [Tversky and Kahneman, 1992] to explain several empirical observations that challenge the standard expected utility theory.
I.6 Connection to General Utility RL and Convex RL in finite trials
$\max_{\pi}F(d_{\rho}^{\pi})\,$, (18)
where $F$ is the real valued utility function defined on the set of probability measures over the state or state-action space, $\rho$ is the initial state distribution and $d_{\rho}^{\pi}$ is the state (or sometimes state-action) occupancy measure induced by the policy $\pi\,.$ This problem captures the standard RL problem as a particular case by considering a linear functional $F$ defined using a fixed given reward function. Recently, motivated by practical concerns, Mutti et al. [2023] argued for the relevance of a variation of the problem under the qualification of convex RL in finite trials. They introduce for this the empirical state distributions $d_{n}\in\Delta(\mathcal{S})$ defined for every state $s\in\mathcal{S}$ by:
$\max_{\pi}\xi_{n}(\pi):=\mathbb{E}[F(d_{n}^{\pi})]\,.$ (20)
Note that $d_{n}^{\pi}$ is a random variable as it is an empirical state distribution. Observe also that $\lim_{n\to\infty}\xi_{n}(\pi)=F(d_{\rho}^{\pi})$ under mild technical conditions (e.g. continuity and boundedness of $F$). This shows the connection between the above final trial convex RL objective and the general utility RL problem (18). The interesting differences between both problem formulations arise for small values of $n$. Of particular interest, both in this paper and in Mutti et al. [2023], is the single trial RL setting where $n=1$.
which is of the form $\xi_{1}(\pi)$, the single-trial RL objective as defined in Mutti et al. [2023]. Indeed, it suffices to write the following to observe it:
I.7 Choice of the utility function
Appendix J More Details about Section 5 and Additional Experiments
@@ I.7 Choice of the utility function
Hardware and execution time. We have run part of the experiments on a laptop with a 13th Gen Intel Core i7-1360P2.20 GHz CPU and 32 GB of RAM, and the remaining simulations (section 5, App. J.8, J.9) on a MacBook Pro M4 with 48 GB of RAM. Experiments can take a few seconds for the simplest ones, to a few minutes for the ones in the main part. More complex ones take a couple of hours, e.g. for electricity management (App. J.7), Finance (App. J.8) and MuJoCo (App. J.9).
@@
Weight functions. The risk-neutral $w_{+}$ function is simply the identity function. As for the definition of other probability distortion functions $w_{+}$ we use for experiments, we define:
@@
J.1 More details about bandit simulations in Section 5
@@
J.2 CPT compared to Distortion Risk Measures
Figure 13: Illustration of the flexibility of CPT compared to the Distortion Risk Measure. Notice how $w_{-}$ is distinct from both $w_{+}$ and $p\mapsto 1-w_{+}(1-p)$.
J.3 Illustration of the need for stochastic policies in CPT-RL
@@
Figure 15: Results of the experiment on non-deterministic policies and batch size influence, over $100$ runs. The black dots are the medians and the shaded area is the interquartile range.
@@
J.4 Markovian vs Non-Markovian Policies for CPT-RL
@@
J.6 Traffic Control over a Grid
We simulate a car agent navigating a city grid, where central roads are faster but risk higher delays (see fig. 20). The agent must balance speed against risk by avoiding the city center. In risk-neutral settings, the agent favors faster routes, while risk-averse policies avoid the risky central roads. Fig. 20 (center) demonstrates that our algorithm successfully adapts to different risk-weighted objectives. We also consider solving CPT-PO with different utility functions: risk-neutral identity utility, risk-averse KT utility, as well as exponential utility function. Fig. 23 show the corresponding different CPT returns observed. The obtained policies differ depending on the utility function. For examples of risk-neutral/averse policies obtained, see Fig. 19(b) in App. J.5.
@@
Figure 21: The probability distortion function $w_{+}$ used for the traffic control experiment.
@@ 2 $\uparrow$ ? $\uparrow$
Figure 23: Returns along the iterations of our PG algorithm for CPT-PO for traffic control. Shaded areas indicate a range of $\pm$ one standard deviation over 20 runs. See App. J for details.
Implementation details. In both cases, the risk-neutral optimal solution (going around the city center) is also a local optimum for the risk-averse objective, and, because it is a shorter path, is easier to stumble upon by chance when exploring the MDP. This means we have to implement special measures to force exploration. The algorithm used as is is prone to get stuck from time to time in local minima on this example. It would seem that our $w$ function, which is aggressively risk-averse, hinders exploration. To mitigate this, we introduce an entropy regularization term that we add to the score function with a decaying regularization weight in the policy gradient found in Theorem 3, see App. J for further details. We incorporate entropy regularization in the policy gradient as follows:
@@
J.8 Trading in Financial Markets
@@
  • State space: We consider a seven dimensional continuous state space. Features are constructed from the raw stock market data as previously explained. State transitions are described using the provided time series.
@@
J.9 Control on MuJoCo Environments